bella2025:

the population of a town has grown by an average of 2,000 people per year over that last 10 years which equation could represent an appropriate linear model of the population. a y=25,000x +2000 b y=2,000x+25,000 c y=-25,000x+2,000 d y= -2,000x+25,000

Alexis1415:

It is B, y=2,000 + 25,000 because, "In slope-intercept form, y = mx + b​: m is the slope or rate of change b is the y-intercept or, in this case, the starting amount If the town is growing by an average of 2,000 people per year, then 2,000 is the rate of change. x is the number of years Based on the answer choices, it appears the starting amount is 25,000 B.) 2,000x + 25,000" Link: https://www.algebrahouse.com/algebra-1-state-test-practice/representing-with-a-linear-model So basically its rate of change is 2,000 because its PER year, so the starting amount is 25,000.

bella2025:

@alexis1415 wrote:
It is B, y=2,000 + 25,000 because, "In slope-intercept form, y = mx + b​: m is the slope or rate of change b is the y-intercept or, in this case, the starting amount If the town is growing by an average of 2,000 people per year, then 2,000 is the rate of change. x is the number of years Based on the answer choices, it appears the starting amount is 25,000 B.) 2,000x + 25,000" Link: https://www.algebrahouse.com/algebra-1-state-test-practice/representing-with-a-linear-model So basically its rate of change is 2,000 because its PER year, so the starting amount is 25,000.
thank you so much I was so confused I thought it was b but I wasnt so sure

Alexis1415:

ofccc <3 glad to help